Suppose there are 100 identical self-service gasoline stations in the city selling the same types of gasoline. Total market demand (QD) and total market supply (Qs) functions for gasoline in the market are as follows:

QD = 60,000 – 25,000P

Qs= 25,000P

where P is expressed in $ per litre.

 

Based on the given information, calculate the equilibrium price and quantity of gasoline in the market. 



1
Expert's answer
2021-12-07T21:05:51-0500

At equilibrium Qd"=" Qs

"60,000-25,000P=25,000P"

"60,000=50,000P"

"P=1.2"

Substituting for P:

"Qd=60,000-25,000(1.2)""=30,000"

At equilibrium

Price"=" $1.2 per litre

Quantity"=" 30,000